OpenStudy (anonymous):

Zach ran t times around the track at school, which is one-fourth mile. He ran a total of seven-eighths mile. Solve for t.

OpenStudy (anonymous):

@bigrob30117

OpenStudy (anonymous):

Well each time he runs around it = 1/4 of a mile....If he ran a total of 7/8 of a mile....he ran... 1/4t = 7/8 Multiply both sides by 4 to isolate t t=78×4 t = ?

OpenStudy (anonymous):

7/8 x 4. My bad.

OpenStudy (anonymous):

i got a little time left before im kicked out of class plz help

OpenStudy (anonymous):

 Do you see how I got this?